# LatentSync: Audio Conditioned Latent Diffusion Models for Lip Sync

### v1

@@ Abstract
We present LatentSync, an end-to-end lip sync framework based on audio conditioned latent diffusion models without any intermediate motion representation, diverging from previous diffusion-based lip sync methods based on pixel space diffusion or two-stage generation. Our framework can leverage the powerful capabilities of Stable Diffusion to directly model complex audio-visual correlations. Additionally, we found that the diffusion-based lip sync methods exhibit inferior temporal consistency due to the inconsistency in the diffusion process across different frames. We propose Temporal REPresentation Alignment (TREPA) to enhance temporal consistency while preserving lip-sync accuracy. TREPA uses temporal representations extracted by large-scale self-supervised video models to align the generated frames with the ground truth frames. Furthermore, we observe the commonly encountered SyncNet convergence issue and conduct comprehensive empirical studies, identifying key factors affecting SyncNet convergence in terms of model architecture, training hyperparameters, and data preprocessing methods. We significantly improve the accuracy of SyncNet from 91% to 94% on the HDTF test set. Since we did not change the overall training framework of SyncNet, our experience can also be applied to other lip sync and audio-driven portrait animation methods that utilize SyncNet. Based on the above innovations, our method outperforms state-of-the-art lip sync methods across various metrics on the HDTF and VoxCeleb2 datasets. Code and models are publicly available at https://github.com/bytedance/LatentSync.
footnotetext: Corresponding authors
@@ 1 Introduction
The task of lip sync involves the generation of accurate lip movements in a talking video based on corresponding speech, while maintaining the head pose and personal identity. This technique has broad applications in numerous practical domains, such as dubbing, virtual avatars, and video conferencing.
@@
Figure 1: Frameworks comparison between previous diffusion-based lip sync methods and our method.
@@
In the field of lip sync, GAN-based methods [18, 36] remain the mainstream approachs. The main issue with GAN-based methods is that they struggle to scale up [24, 44] to large and diverse datasets due to the unstable training and mode collapse [47, 6]. Recent studies proposed diffusion-based methods [34, 2, 64, 57, 29] for lip sync, allowing the model to easily generalize across different individuals without the need for further fine-tuning on specific identities. However, these methods still have some limitations. Specifically, [34, 2] performs the diffusion process in the pixel space (Fig. 1 a), which restricts its ability to generate high-resolution videos due to the prohibitive hardware requirements. Methods like [64, 57] adopt a two-stage approach: the first stage generates lip motions from audio, and the second stage synthesizes the visual appearance conditioned on the motion (Fig. 1 b). The issue with this two-stage approach is that subtly different sounds may map to the same motion representation, which can result in the loss of nuanced expressions linked to the emotional tone of the speech.
To solve the above problems, we propose LatentSync, a novel end-to-end lipsync framework based on audio conditioned latent diffusion models [41], without the need of any intermediate 3D representation [3] or 2D landmarks [5, 30], as shown in Fig. 1 (c). LatentSync leverages the powerful generative capabilities of Stable Diffusion [41] to directly capture the intricate audio-visual correlations, facilitating the generation of dynamic and lifelike talking videos.
@@
Furthermore, we found that the lip-sync accuracy of our model was quite low without SyncNet [9, 36] supervision. We explored two methods to add SyncNet supervision to latent diffusion models [41]: (a) adding SyncNet supervision after decoding to pixel space, and (b) training a SyncNet in the latent space. However, we discovered that SyncNet has convergence issues in both high-resolution pixel space and latent space. To solve this problem, we conducted comprehensive empirical studies in the aspects of model architecture, training hyperparameters, and data preprocessing methods, allowing the SyncNet to converge easily. We achieved the unprecedented 94% accuracy on HDTF [62].
Additionally, we found that in comparison to single-step lip sync methods [36, 34, 61, 18], diffusion-based lip sync approaches [34, 2], which employ multi-step generation processes, exhibit inferior temporal consistency. We hypothesize that this might be due to the inconsistency in the diffusion process across different frames. Inspired by [58, 15], we proposed a new method called TREPA, which uses a large-scale self-supervised video model VideoMAE-v2 [52] to extract a powerful representation rich in temporal information. The distance between the temporal representations of the generated consecutive frames and the ground truth consecutive frames is used as an extra loss. This method is very similar to LPIPS [59], with the key difference being that this loss leverages temporal representations. We found that this simple method effectively improves the temporal consistency while preserving the lip-sync accuracy of our model.
In summary, we make the following contributions: (1) We proposed an end-to-end lip sync framework based on audio conditioned latent diffusion model without any intermediate motion representation. (2) To solve the convergence issue of SyncNet, we conducted empirial studies in the aspects of model architecture, training hyperparameters, data preprocessing methods and identified the key factors affecting SyncNet convergence. (3) We proposed the TREPA to improve temporal consistency while preserving the lip-sync accuracy.
@@ 2.1 Diffusion-based Lip Sync
[34] and [2] are both end-to-end lip sync methods based on pixel space audio conditioned diffusion models. [57] uses diffusion models in the first stage to complete audio-to-motion conversion and uses a VAE [26] in the second stage for motion-to-image generation. [64] uses transformers in the first stage to convert audio to motion and employs diffusion models in the second stage for motion-to-image generation. [29] uses diffusion autoencoders [37] to convert the masked images into semantic latent codes in the first stage, and uses diffusion models to generate an image conditioned on the semantic latent codes and audio in the second stage, which also belongs to pixel space diffusion models.
@@ 2.2 Non-diffusion-based Lip Sync
Wav2Lip [36] is the most classic lip sync method that introduced using a pretrained SyncNet [9] to supervise the training of the generator. [20] trains a VQ-VAE [12] to encode faces and head poses, and then trains the lip sync generator in the quantized space to generate high-resolution images. StyleSync [18] follows the overall framework of Wav2Lip, with its main innovation being the use of StyleGAN2 [25] as the generator backbone. [8] divides lip sync into three components: semantic-guided reenactment network, lip sync network, and identity-aware refinement and enhancement. DINet [63] deforms the feature maps to generate mouth shapes conditioned on the driving audio. MuseTalk [61] uses the architecture of Stable Diffusion [41] for inpainting, but it does not perform diffusion process and uses a discriminator for adversarial learning [17], making it more like a GAN-based framework.
@@ 3 Method
Figure 2: The overview of our LatentSync framework. We use the Whisper [39] to convert melspectrogram into audio embeddings, which are then integrated into the U-Net [42] via cross-attention layers. The reference and masked frames are channel-wise concatenated with noised latents as the input of U-Net. In the training process, we use one-step method to get estimated clean latents from predicted noises, which are then decoded to obtain the estimated clean frames. The TREPA, LPIPS [59] and SyncNet loss [36] are added in the pixel space.
@@ 3.1 LatentSync Framework
The overview of the LatentSync framework is shown in Fig. 2. Our backbone network generates the video frame by frame without temporal modeling between frames. The temporal consistency during the generation process relies entirely on the audio window. This frame-by-frame generation approach is beneficial for occlusion strategies in dubbing scenarios, as we can easily skip frames where the mouth is occluded and lip sync is not needed. Our framework is based on image-to-image inpainting [43], so it requires a masked image as input. Additionally, to integrate the visual features of the face from the original video, we need to input reference images. The reference images are randomly selected during training to avoid learning shortcuts [16], and they are taken from the current frames during inference. We concatenate these inputs along the channel dimension, so the total input consists of 13 channels (4 channels for noise latent, 1 channel for the mask, 4 channels for the masked image, and 4 channels for the reference image). At the beginning of training, we initialized our model with the parameters of SD 1.5 [41], except for the first conv_in layer with 13 channels and cross-attention layers of dimension 384, which are randomly initialized. Theoretically, our framework is compatible with more advanced latent diffusion models such as SDXL [35], SD3 [13], and FLUX [28], but we only conducted experiments on SD 1.5 due to hardware limitations.
@@
SyncNet supervision. Our model predicts in the noise space, while SyncNet [9, 36] requires an input in the image space. To address this issue, we use the predicted noise $\epsilon_{\theta}(z_{t})$ to obtain the estimated $\hat{z}_{0}$ in one step, which can be formulated as:
@@
Another problem is that our model makes predictions in the latent space. We explored two methods to add SyncNet supervision to latent diffusion models: (a) Decoded pixel space supervision, which trains SyncNet in the same way as [36]. (b) Latent space supervision, which requires training a SyncNet in the latent space. The visual encoder input of this SyncNet is the latent vectors obtained by the VAE [12, 26] encoding. The illustration is shown in Fig. 3. We found that training SyncNet in the latent space converges worse compared to training it in the pixel space. We speculate that this is due to information loss in the lip area during the VAE encoding process. The poorer convergence of the latent space SyncNet also results in the worse audio-lip sync accuracy for the supervised diffusion models, according to the experimental results in Sec. 4.3. Therefore, we use decoded pixel space supervision in the LatentSync framework.
@@
Two-stage training. In the first stage, we do not decode to the pixel space and do not add SyncNet loss. The purpose of the first stage is to allow the U-Net to learn visual features using a large batch size. Therefore, the training objective of the first stage has only a simple loss [23]:
@@
In the second stage of training, we used the decoded pixel space supervision method to add SyncNet loss. Compared to [36], we increased the input frame length of SyncNet to 16. For detailed changes to the SyncNet structure, please refer to Sec. 3.3. By using the 16 decoded video frames $\mathcal{D}(\hat{z}_{0})_{f:f+16}$ and the corresponding audio sequence $a_{f:f+16}$, the SyncNet loss can be formulated as:
@@
We identified the optimal coefficients through empirical analysis as follows: $(\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4})=(1,0.05,0.1,10)$
Mixed noise model. To ensure that the model correctly learns temporal information, the input noises also need to have temporal consistency; otherwise the model will fail to learn properly. The simplest method is applying the same noise to all frames [34]. We further found that diffusion models trained with the mixed noise model [14] exhibit better temporal consistency. The method generates two distinct noise vectors, $\epsilon_{\text{shared}}$ and $\epsilon_{\text{ind}}$. $\epsilon_{\text{shared}}$ represents a common noise vector applied across all 16 video frames, whereas $\epsilon_{\text{ind}}$ corresponds to the frame-specific noise. The final noise is obtained through a linear combination of these two vectors.
$\epsilon^{f}=\epsilon_{\text{shared}}+\epsilon_{\text{ind}}^{f}$ (6)
We used the mixed noise model in both two training stages, and we applied the same initial latent noise to all frames in the video during inference stage.
Affine transformation. During the data preprocessing stage, affine transformation was employed to achieve face frontalization. This approach [18] helps the model to effectively learn facial features particularly in challenging scenarios such as side-profile views. We applied a mask that covers the entire face to prevent information leakage, ensuring that the model does not infer lip movements based on facial movements. The position and shape of the mask are fixed. We do not use detected landmarks [30, 5] to draw the mask as they will leak information about facial movements. The affine transformation and fixed mask are illustrated in Fig. 4.
Figure 4: The illustration of affine transformation and fixed mask.
@@ 3.2 Temporal Representation Alignment
TREPA aligns the temporal representations of the generated image sequences with those of ground truth image sequences. The insight behind this method is that merely employing distance loss between individual images only improves the content quality of single generated images but does not enhance the temporal consistency of the generated image sequence. In contrast, temporal representations capture temporal correlation within image sequences, enabling the model to focus on improving the overall temporal consistency. We employed a large-scale self-supervised video model VideoMAE-v2 [52] to extract temporal representations. Due to its unsupervised training on large-scale unlabeled datasets, the model’s temporal representations exhibit strong generalization capabilities, robustness, and high information density.
@@
Observations. The temporal layer [19, 4] adds temporal self-attention layers to the U-Net to enhance temporal consistency. We initially attempted to apply the temporal layer to LatentSync, but found that it significantly compromised lip-sync accuracy. In contrast, our TREPA not only avoids harming lip-sync accuracy but even improves it to some extent. We hypothesize this is because the temporal layer adds additional parameters to the U-Net, causing some of the gradients in backpropagation to be distributed to the temporal layer parameters, weakening the learning of the audio cross-attention layer parameters. On the other hand, TREPA does not add extra parameters; to improve temporal consistency, the model must rely more effectively on the information in the audio window (since we know that lip-sync models use the audio window to capture temporal information). As a result, the audio cross-attention layers are further trained and strengthened. Refer to the quantitative results in Sec. 4.3.
Another interesting finding is that TREPA can also be applied to improve the temporal consistency and lip-sync accuracy of Wav2Lip [36], according to the experimental results in Sec. 4.3. Theoretically, any video generation model that generates multiple consecutive frames at once can utilize TREPA to enhance the temporal consistency.
3.3 Empirical Studies on SyncNet Convergence
Many researchers have reported the convergence problem of SyncNet on GitHub issues [32, 56, 40, 33], a typical characteristic of this issue is that the training loss gets stuck at 0.69 and fails to decrease further. However, no academic work has provided a detailed analysis of this problem to date. In this section, we analyze the SyncNet convergence problem and identify several critical factors affecting the convergence of SyncNet through various ablation studies. Importantly, we preserved SyncNet’s original training framework and employed the same contrastive loss function as the original paper [9]. Therefore, our experience can be applied to many lip sync [36, 18, 34] and audio-driven portrait animation methods [31] that utilize SyncNet.
Why is the SyncNet training loss stuck at 0.69? According to the classic framework of SyncNet [36, 9], we randomly provide SyncNet with positive and negative samples with a 50% probability during the training stage. The output of SyncNet is a probability distribution of whether the sample is positive or negative. We define $p(x=1)$ as the probability that the sample is positive, and $p(x=0)$ as the probability that the sample is negative. Let $p(x)$ be the true probability distribution and $q(x)$ the predicted probability distribution. When $q(x=1)\approx q(x=0)\approx 0.5$ and batch size $N$ is sufficiently large, the training loss of SyncNet is (proof in Appendix B):

### v2

# LatentSync: Taming Audio-Conditioned Latent Diffusion Models for Lip Sync with SyncNet Supervision

@@ Abstract
End-to-end audio-conditioned latent diffusion models (LDMs) have been widely adopted for audio-driven portrait animation, demonstrating their effectiveness in generating lifelike and high-resolution talking videos. However, direct application of audio-conditioned LDMs to lip-synchronization (lip-sync) tasks results in suboptimal lip-sync accuracy. Through an in-depth analysis, we identified the underlying cause as the “shortcut learning problem”, wherein the model predominantly learns visual-visual shortcuts while neglecting the critical audio-visual correlations. To address this issue, we explored different approaches for integrating SyncNet supervision into audio-conditioned LDMs to explicitly enforce the learning of audio-visual correlations. Since the performance of SyncNet directly influences the lip-sync accuracy of the supervised model, the training of a well-converged SyncNet becomes crucial. We conducted the first comprehensive empirical studies to identify key factors affecting SyncNet convergence. Based on our analysis, we introduce StableSyncNet, with an architecture designed for stable convergence. Our StableSyncNet achieved a significant improvement in accuracy, increasing from 91% to 94% on the HDTF test set. Additionally, we introduce a novel Temporal Representation Alignment (TREPA) mechanism to enhance temporal consistency in the generated videos. Experimental results show that our method surpasses state-of-the-art lip-sync approaches across various evaluation metrics on the HDTF and VoxCeleb2 datasets. Code and models are publicly available at https://github.com/bytedance/LatentSync.
@@ 1 Introduction
The lip sync [29, 28, 14, 52] is a video editing task, which regenerates the lip movements of a talking person according to the given audio, while maintaining the head pose and personal identity. This technique has broad applications in numerous practical domains, such as visual dubbing, virtual avatars, and video conferencing.
@@
Figure 1: Frameworks comparison between previous diffusion-based lip-sync methods and our method.
@@
In the field of lip sync, GAN-based methods [14, 29] remain the mainstream approachs. The main issue with these methods is that they struggle to scale up [21, 36] to large and diverse datasets due to the unstable training [2, 35] and mode collapse [39, 6]. Recent studies proposed diffusion-based methods [28, 3, 53, 48, 25] for lip sync, allowing the model to easily generalize across different individuals without the need for further fine-tuning on specific identities. However, these methods still have some limitations. Specifically, [28, 3] perform the diffusion process in the pixel space (Fig. 1 a), which restricts its ability to generate high-resolution videos due to the prohibitive hardware requirements. Other methods [53, 48] adopt a two-stage approach: the first stage generates lip motions from audio, and the second stage synthesizes the visual appearance conditioned on the motion (Fig. 1 b). The issue with this two-stage approach is that subtly different sounds may map to the same motion representation, leading to the loss of nuanced expressions linked to the emotional tone of the speech.
To address the above limitations, we propose LatentSync, an end-to-end lip sync framework based on audio-conditioned LDMs [33] to generate lifelike and high-resolution talking videos, as shown in Fig. 1 (c). We initially tried directly applying methods from the field of audio-driven portrait animation, such as EMO [41] and Hallo [46]. However, the results showed poor lip-sync accuracy. We delved deeper into this phenomenon and identified the “shortcut learning problem” [12] inherent in lip sync task.
The shortcut learning problem in lip sync. Lip-sync methods are typically based on a video-to-video inpainting framework, the model receives masked frames and audio as inputs. Unexpectedly, the audio-conditioned LDMs tends to predict lip movements based on visual information around the lips, such as facial muscles, eyes, and cheeks, while ignoring the audio information. We conducted an experiment to validate the existence of the shortcut learning problem and the effectiveness of SyncNet supervision [29] in mitigating this issue. We trained the audio-conditioned LDMs with masks of different sizes, both with and without SyncNet supervision. We used the sync confidence score [9] to evaluate the synchronization accuracy between audio and lip movements.
Figure 2: The shortcut learning problem in the lip-sync task. Higher sync confidence score means better lip-sync accuracy.
As shown in Fig. 2, without SyncNet supervision, as the mask size increases, the visual information available in the masked frame for inferring lip movements decreases, forcing the model to rely more on audio information, thereby improving lip-sync accuracy, and vice versa. In contrast, with SyncNet supervision, the model remains focused on learning audio-visual correlations across different mask sizes.
Why is SyncNet supervision not necessary for audio-driven portrait animation methods? These methods [41, 46, 7, 20] are based on image-to-video framework. Since they do not involve input masked frames, they do not suffer from the shortcut learning problem.
@@
The previous work Diff2Lip [28] has explored how to add SyncNet supervision to pixel-space diffusion models. However, how to effectively apply SyncNet to latent diffusion models [33] remains unclear. Specifically, we explored two methods to incorporate SyncNet supervision into latent diffusion models: (a) Decoded pixel space supervision and (b) Latent space supervision. Furthermore, We found that SyncNet struggles to converge in both latent space and high-resolution pixel space. Since the convergence of SyncNet significantly impacts its supervision effectiveness, and ultimately affects the lip-sync accuracy of the supervised model, the convergence issue of SyncNet is highly valuable for research. Therefore, we conducted the first comprehensive empirical studies in the aspects of model architecture design, training hyperparameters, and data preprocessing techniques, introducing the StableSyncNet with an architecture designed for stable convergence. Our StableSyncNet achieved the unprecedented 94% accuracy on HDTF [51].
Additionally, we observed that high-frequency details in the generated talking videos, such as teeth, lips, and facial hair, exhibit flickering artifacts. We propose TREPA, a novel method designed to enhance temporal consistency and reduce such artifacts.
In summary, we made the following contributions: (1) We proposed LatentSync, the first lip-sync method that utilizes audio-conditioned LDMs to achieve end-to-end lifelike lip sync on high-resolution videos, incorporating TREPA to enhance the temporal consistency in the generated videos. (2) We identified the shortcut learning problem in lip-sync task and explored different methods to incorporate SyncNet supervision into audio-condtioned LDMs. (3) We conducted the first comprehensive empirical studies to identify key factors affecting SyncNet convergence, introducing the StableSyncNet for stable convergence.
Figure 3: The overview of our LatentSync framework. We use the Whisper [32] to convert melspectrogram into audio embeddings, which are then integrated into the U-Net [34] via cross-attention layers. The reference and masked frames are channel-wise concatenated with noised latents as the input of U-Net. In the training process, we use a one-step method to get estimated clean latents from predicted noises, which are then decoded to obtain the estimated clean frames. The TREPA, LPIPS [49] and SyncNet loss [29] are added in the pixel space.
@@ 2.1 Diffusion-based Lip Sync
Diff2Lip [28] and DrivenVideoEditing [3] are both end-to-end lip sync methods based on pixel-space audio-conditioned diffusion models. MyTalk [48] uses diffusion models in the first stage to complete audio-to-motion conversion and uses a VAE [23] in the second stage for motion-to-image generation. StyleSync [53] uses transformers in the first stage to convert audio to motion and employs diffusion models in the second stage for motion-to-image generation. DiffDub [25] uses diffusion autoencoders [30] to convert the masked images into semantic latent codes in the first stage, and uses diffusion models to generate an image conditioned on the semantic latent codes and audio in the second stage.
@@ 2.2 Non-diffusion-based Lip Sync
Wav2Lip [29] is the most classic lip sync method that introduced using a pretrained SyncNet [9] to supervise the training of the lip-sync generator. [16] trains a VQ-VAE [11] to encode faces and head poses, and then trains the lip sync generator in the quantized space to generate high-resolution images. StyleSync [14] follows the overall framework of Wav2Lip, with its main innovation being the use of StyleGAN2 [22] as the generator backbone. VideoReTalking [8] divides lip sync into three components: semantic-guided reenactment network, lip sync network, and identity-aware refinement and enhancement. DINet [52] deforms the feature maps to generate mouth shapes conditioned on the driving audio. MuseTalk [50] uses the architecture of Stable Diffusion [33] for inpainting, but it does not perform diffusion process and uses a discriminator for adversarial learning [13], making it more like a GAN-based framework.
2.3 Audio-driven Portrait Animation
Many people may confuse lip sync with audio-driven portrait animation. These two tasks have some similarities but are actually completely different tasks. Lip sync is based on video-to-video editing framework, which needs to keep areas other than the mouth the same as in the input video. Audio-driven portrait animation is based on image-to-video animation framework, which can change the head movement and even facial expressions. Although there are already some audio-driven portrait animation methods based on audio-conditioned LDMs, such as [41, 7, 46, 20], they cannot be directly applied to the lip sync task due to the shortcut learning problem.
@@ 3.1 LatentSync Framework
The overview of the LatentSync framework is shown in Fig. 3. The framework is based on video-to-video inpainting with the temporal modeling by temporal layer [15]. To incorporate the visual features of the face from the input video, reference frames are introduced as additional inputs. During training, these frames are randomly selected, while during inference, they are taken from the current frames. We concatenate different inputs along the channel dimension, making the total input of U-Net to be 13 channels (4 channels for noise latent, 1 channel for the mask, 4 channels for the masked frame, and 4 channels for the reference frame). At the beginning of training, the model is initialized with the parameters of SD 1.5 [33], except for the first conv_in layer with 13 channels and cross-attention layers of dimension 384, which are randomly initialized.
@@
Affine transformation and fixed mask. During the data preprocessing stage, affine transformation was employed to perform face frontalization. This approach [14] helps the model to effectively learn facial features particularly in challenging scenarios such as side-profile views. We applied a mask that covers the entire face to minimize the model’s tendency to learn visual-visual shortcuts. The position and shape of the mask are fixed. We do not use the detected landmarks [26, 5] to draw the mask, as moving landmarks also provide cues about lip movements. The affine transformation and fixed mask are illustrated in Fig. 4.
Figure 4: The illustration of affine transformation and fixed mask.
SyncNet supervision. Latent diffusion models predict in the noise space, while SyncNet [9, 29] requires an input in the image space. To address this issue, we use the predicted noise $\epsilon_{\theta}(z_{t})$ to obtain the estimated $\hat{z}_{0}$ in one step, which can be formulated as:
@@
Another problem is that latent diffusion models make predictions in the latent space. We explored two methods to incorporate SyncNet supervision into latent diffusion models: (a) Decoded pixel space supervision, which trains SyncNet in the same way as Wav2Lip [29]. (b) Latent space supervision, which requires training a SyncNet in the latent space. The visual encoder input of this SyncNet is the latent vectors obtained by the VAE [11, 23] encoding. The illustration is shown in Fig. 5. Our empirical analysis in Sec. 5.3 reveals that training SyncNet in the latent space exhibits inferior convergence compared to training in the pixel space. This degradation may arise from information loss in the lip region during the VAE encoding process. The poorer convergence of SyncNet in the latent space adversely impacts the lip-sync accuracy of the supervised diffusion models, according to the experimental results in Sec. 5.3. Therefore, we finally choose the decoded pixel space supervision in the LatentSync framework.
@@
3.2 Two-Stage Training Strategy
The decoded pixel space supervision has a problem that the activations in VAE decoding need to be stored for backpropagation, which significantly increases GPU memory consumption. To mitigate this issue, we designed a two-stage training strategy: in the first stage, the model learns to extract features from reference frames and develops inpainting capabilities, which we refer to as learning visual features. In the second stage, it learns audio-visual correlations under SyncNet supervision. This approach eliminates the need for the VAE decoding process in the first stage, enabling model to learn visual features with a larger batch size. We observed that audio-conditioned LDMs typically spend more time learning visual features and less time learning audio-visual correlations. Therefore, this two-stage training strategy allows the model to efficiently learn both visual features and audio-visual correlations. We provide the formal definitions of training objectives in the following paragraphs.
In the first stage of training, we do not add the temporal layer [15, 4] and train all parameters of the U-Net. The training objective has only a simple loss [19]:
@@ 3.2 Two-Stage Training Strategy
In the second stage of training, we only train the temporal layer and audio layer while freezing the other parameters of the U-Net. Suppose we have 16 decoded video frames $\mathcal{D}(\hat{z}_{0})_{f:f+16}$ and the corresponding audio sequence $a_{f:f+16}$, the SyncNet loss can be formulated as:
@@ 3.3 Temporal Representation Alignment
TREPA aligns the temporal representations of the generated image sequences with those of ground truth image sequences. The insight behind this method is that merely employing distance loss between individual images improves the content quality of single generated images but does not enhance the temporal consistency of the generated image sequence. In contrast, temporal representations capture temporal correlation within image sequences, enabling the model to focus on improving the overall temporal consistency. We employed a large-scale self-supervised video model VideoMAE-v2 [44] to extract temporal representations. Due to its unsupervised training on large-scale unlabeled datasets, the model’s temporal representations exhibit strong generalization capabilities, robustness, and high information density.
@@
4 Empirical Studies on SyncNet Convergence
Our experiments reveal that SyncNet is difficult to converge in both latent space and high-resolution pixel space, a typical characteristic of this issue is that the training loss gets stuck at 0.69 and fails to decrease further. In this section, we analyze the SyncNet convergence problem and identify several critical factors affecting the convergence through various ablation studies. Importantly, we preserved SyncNet’s original training framework and employed the same contrastive loss function as Wav2Lip [29]. Therefore, our experience can be applied to many lip-sync [29, 14, 28, 8, 52, 50] and audio-driven portrait animation methods [27, 47] that utilize SyncNet.
Why is the SyncNet training loss stuck at 0.69? According to the classic training framework of SyncNet [29, 9], we randomly provide SyncNet with positive and negative samples with a 50% probability during the training stage. The output of SyncNet is a probability distribution of whether the sample is positive or negative. We define $p(x=1)$ as the probability that the sample is positive, and $p(x=0)$ as the probability that the sample is negative. Let $p(x)$ be the true probability distribution and $q(x)$ the predicted probability distribution. When $q(x=1)\approx q(x=0)\approx 0.5$ and batch size $N$ is sufficiently large, the training loss of SyncNet is (proof in Appendix A):